### v5

# Role of magnetic ions in the thermal Hall effect of the paramagnetic insulator TmVO4

Abstract
In a growing number of materials, phonons have been found to generate a thermal Hall effect, but the underlying mechanism remains unclear. Inspired by previous studies that revealed the importance of Tb3+ ions in generating the thermal Hall effect in a family of pyrochlores, we investigated the role of Tm3+ ions in TmVO4, a paramagnetic insulator with a different crystal structure. We observe a negative thermal Hall conductivity in TmVO4 with a magnitude such that the Hall angle, /, is approximately 1 10-3 at = 15 T and = 20 K, typical for a phonon-generated thermal Hall effect. In contrast to the negligible found in the nonmagnetic pyrochlore analog (where the Tb3+ ions are replaced with Y3+), we observe a negative in YVO4 with a Hall angle of magnitude comparable to that of TmVO4. This shows that the Tm3+ ions are not essential for the thermal Hall effect in this family of materials. Interestingly, at an intermediate Y concentration of = 0.3 in Tm1-xYxVO4, was found to have a positive sign, pointing to the importance of impurities in the thermal Hall effect of phonons.
pacs:
Valid PACS appear hereI INTRODUCTION
In the last decade, a thermal Hall effect has been measured in a number of insulators where phonons are the main heat carriers, including multiferroics such as Fe2Mo3O8 [1], cuprates such as La2CuO4 [2, 3] and Nd2CuO4 [4], nonmagnetic SrTiO3 [5], and the antiferromagnet Cu3TeO6 [6], see Fig. 1. In Tb3Ga5O12 [7], the first insulator in which a thermal Hall signal was detected, the effect was attributed to skew scattering of phonons by superstoichiometric Tb3+ impurities [8].
Since then, a number of theoretical scenarios have been proposed to explain the origin of the phonon thermal Hall effect. Some attribute the thermal Hall conductivity to the Berry curvature of phonon bands [9]. Others link it to various types of spin-lattice coupling [10, 11, 12, 13]. In yet others, the role of impurities is considered important [14, 15, 16, 17], but it remains unclear which of these scenarios applies to what material.
In a previous study on pyrochlores [18], a sizable was observed in Tb2Ti2O7 but a negligible was detected in Y2Ti2O7. Although the signal in Tb2Ti2O7 was originally attributed to some exotic neutral excitations linked to the spin-liquid nature of the system [18], it was later argued that phonons are, in fact, responsible for the thermal Hall effect in this material [19], since a signal of comparable magnitude is still observed when 70% of the Tb3+ ions are replaced by Y3+ ions, and the spin state of the system is profoundly altered. Irrespective of the underlying mechanism, this study does show that Tb3+ ions are essential for generating a thermal Hall effect in these pyrochlores.
Here we report a similar study carried out on a different oxide, TmVO4, wherein we investigate what happens to the thermal Hall effect when we substitute the magnetic ion Tm3+ for the nonmagnetic ion Y3+.
The insulator TmVO4 exhibits a rich set of phenomena at low temperatures. It undergoes a cooperative Jahn-Teller phase transition at K in zero field, and at T at (for ) [25]. (Note that we use as a shorthand notation for .) The ordered state consists of a simultaneous ferroquadrupole order of the local 4 electronic orbitals of each Tm atom, accompanied by a structural transition, by which the crystal structure goes from tetragonal at high temperature (41/ space group [21]) to orthorhombic at low temperature [26, 27, 25, 28, 29]. However, no magnetically ordered state has been reported in TmVO4 [25]. Our own study will focus on temperatures above .
II METHODS
II.1 Samples
The following three rare-earth vanadates were studied: TmVO4, YVO4, and Tm0.7Y0.3VO4. The samples were grown using a flux growth method [30, 31]. They are needle-shaped with their long axis along the axis, which is the easy axis of magnetization [32]. Because TmVO4 has a large anisotropy in its -factor ( = 10 and = 0) [33], even a slight misalignment of the magnetic field can cause a large torque on the sample, which can detach it from its mount.
Two measures were adopted to prevent this. First, the thermal transport contacts on TmVO4 and Tm0.7Y0.3VO4 samples were made with silver epoxy to ensure strong contacts that would adhere to a very smooth surface and would not peel off. These contacts were solidified for about 10 minutes on a hot plate at 150∘ C. Note that the contacts on the YVO4 samples were done with silver paint, since there is no magnetic torque in this case. For the heater contact, a 50-m diameter silver wire was used and for the thermal transport contacts, wires with a diameter of 25 m were used.
In addition, a small wood stick was positioned along the TmVO4 and Tm0.7Y0.3VO4 samples, and was linked to the sample by the heater wire. The thermal conductivity was measured (in zero field) with and without the wood stick and only a 5 % (8 %) difference was observed at the peak temperature for TmVO4 (Tm0.7Y0.3VO4). The sample dimensions (contact dimensions) are given in Table 1.
II.2 Thermal transport measurements
To perform a conventional thermal transport study, a five-contact measurement is done using a steady-state method at a fixed field, as sketched in Fig. 2. A heat current () is generated along the axis from the heater contact on one end of the sample and the copper block on the other end. The heater is a strain gauge of 5 k whose resistance does not change with temperature or field. (No potential strain is conferred to the sample since the heater is glued with GE varnish to a silver pad.) This heat current was applied along the axis of the sample (the axis in Fig. 2), while the field was applied along the axis (the axis in Fig. 2). Note that the field ( = 10 or 15 T) was applied at high temperatures (at 83 K) and the samples were cooled down to K using a variable temperature insert (VTI). This procedure ensures that no hysteresis due to the sample is generated.
| (mm) | (mm) | (mm) | |
| TmVO4 | 0.66 | 0.24 | 0.04 |
| YVO4 | 0.49 | 0.25 | 0.06 |
| Tm0.7Y0.3VO4 | 0.65 | 0.45 | 0.04 |
Temperature steps from K to 80 K were done at fixed fields and all the temperature signals were stable before measuring each temperature difference ( and in Fig. 2) using type-E (chromium/constantan) thermocouples. Note that a test has been done previously to show that thermocouples and chip-based thermometers (Cernox) yield the same thermal Hall data (see Appendix C in ref. [34]). Also, note that for all the voltages, the background voltage (when the heat is off) is subtracted from the measured signal (when the heat is on).
The conductivities and were measured as described elsewhere [2, 3, 4]. Note that the transverse temperature difference is obtained by measuring in both field polarities ( and ). We checked that our technique is independent of the particular choice of sequence, i.e., whether we measure first and then or if we measure instead first and then . This means that our experimental setup and our procedure do not generate any (technique dependent) hysteresis. Then, we antisymmetrize the measured signals to get rid of any longitudinal contribution due to a possible slight misalignment of the transverse contacts:
The sign of is determined by measuring a sample of known with the same setup and wiring. In a metallic sample with Hall coefficient 0 (thus an electrical Hall conductivity ), the thermal Hall conductivity will be positive (at least in the = 0 limit) since both thermal and electrical conductivities are related by the Wiedemann-Franz law : for 0.
The thermal conductivity is defined as:
where is the heat current (in W) and is a geometric factor (; see Table 1 for sample dimensions). The thermal Hall conductivity is defined as:
Note that here we assume that (i.e., that ), but this is not quite right since the axis conductivity is not identical to the axis conductivity. This implies that the amplitude of reported here is not quite accurate, and would need to be multiplied by the anisotropy factor . Given the needle shape of our samples, measurements with have not been done.
III RESULTS
III.1 Thermal conductivity,
The thermal conductivity of our three samples is displayed in Fig. 3. All three curves are typical of insulators, with a peak near 20-30 K.
The thermal conductivity of TmVO4 was measured previously [35]. These early measurements were done with (along the easy axis [32]) for T and for K, whereas our measurements were done with and , covering a larger temperature and field range. This previous study reports a slightly higher peak, perhaps due to larger samples and higher sample quality. The field was found to increase at low temperatures ( K). In our study, we also see an increase in with at low temperatures [ K; inset of Fig. 4(a)].
The zero-field curves are compared in Fig. 3(d), where we see that in YVO4 is much larger than in the other two samples. The magnitude of in the stoichiometric sample of paramagnetic TmVO4 is much lower presumably because phonons scatter on the crystal field levels of the Tm3+ ions. This is also presumed to be the case in the frustrated spin system Tb2Ti2O7, where the scattering of phonons would involve the crystal field levels of the Tb3+ ions, as argued previously [8]. This scattering process is much stronger for Tb3+ than Tm3+, resulting in a value of W/mK at K in Tb2Ti2O7 [36], compared to W/mK in TmVO4 [Fig. 3(a)].
As in the case of Tb2Ti2O7, where the field dependence of is exceptionally strong [36], we view the sizable field dependence of in TmVO4 [Fig. 3(a)] and Tm0.7Y0.3VO4 [Fig. 3(c)], as a confirmation that phonons are scattered by the crystal field levels of the Tm3+ ions. The weaker field dependence seen in YVO4 [Fig. 3(b)] is consistent with that picture. We attribute the fact that in Tm0.7Y0.3VO4 is even smaller than in TmVO4 to the extra disorder introduced by the Y impurities [Fig. 3(d)].
In Fig. 4, we plot versus , to emphasize the data at low temperature. We see that below 10 K or so, the field dependence of in TmVO4 becomes very strong and opposite in sign relative to its behavior at higher temperature [inset of Fig. 4(a)]. We speculate that this change of behavior is associated with the quadratic splitting of the crystal field levels of the Tm3+ ions with field (second-order Zeeman interaction for //a) [25] which changes the resonant phonon scattering (for phonons of appropriate symmetry).
Interestingly, this change in field dependence, as the temperature is raised, was actually captured by theoretical calculations that consider the main scattering mechanism of phonons to be resonant scattering on electronic levels of the Tm3+ ions [37, 38], albeit for temperatures much closer to the transition.
III.2 Thermal Hall conductivity,
The thermal Hall conductivity of our three samples is displayed in Fig. 5. Only data at T are shown, but data at T were also taken; they are similar but smaller in magnitude, roughly in proportion to the field amplitude. The first observation is that all three samples exhibit a non-negligible thermal Hall conductivity. The surprise is that the sign of is negative in the two stoichiometric materials, TmVO4 and YVO4, but it is positive in the disordered sample Tm0.7Y0.3VO4.
We see that the temperature at which peaks (Fig. 5) is roughly the same as the temperature at which the phonon-dominated peaks (Fig. 3), for example, K in TmVO4 and K in YVO4. As argued before for other materials [5, 4, 6, 24], this is an argument in support of phonons being the heat carriers responsible for the thermal Hall signal in these materials.
Looking at Fig. 5, we also notice that the magnitude of appears to roughly scale with the magnitude of seen in Fig. 3(d). This is confirmed when looking at the thermal Hall angle, plotted as versus in Fig. 6. Indeed, we see that at K in both YVO4 and Tm0.7Y0.3VO4, even though the magnitude of is 25 times larger in YVO4 (Fig. 5). This is because is also roughly 25 times larger in YVO4 (Fig. 3).
IV DISCUSSION
In the pyrochlore study, a non-zero (and positive) thermal Hall effect was measured in Tb2Ti2O7, but a negligible was detected in Y2Ti2O7 [18]. It was initially speculated that the thermal Hall signal in Tb2Ti2O7 was due to some exotic spin excitations [18]. However, a subsequent study found that the signal remains as strong when 70% of the magnetic Tb3+ ions are replaced by nonmagnetic Y3+ ions, thereby dramatically altering the frustrated spin-liquid state of pure Tb2Ti2O7 [19]. This lead to the conclusion that phonons are, in fact, the heat carriers responsible for the thermal Hall effect in these pyrochlores. Moreover, the fact that becomes negligible in Y2Ti2O7, when all Tb is replaced by Y, suggests that magnetism plays a key role in the generation of the phonon thermal Hall effect in these pyrochlores.
Turning to our own study, we also observe a non-zero in TmVO4, another oxide with magnetic ions (Tm3+), and we also attribute this Hall effect to phonons, as argued above. In our comparison with Tb2Ti2O7, we see two differences. The first is the fact that our nonmagnetic parent compound, YVO4, also displays a non-zero , unlike the negligible signal of Y2Ti2O7. So in these vanadates, the phonon Hall effect does not depend crucially on the magnetic moment associated with Tm3+ ions. Note that a non-zero thermal Hall effect from phonons has been observed in other nonmagnetic insulators, such as strontium titanate [5] and black phosphorous [24].
The second difference is the sign of : positive in Tb2Ti2O7, negative in TmVO4. The sign of the phonon Hall conductivity is an entirely open question, on which existing theories shed little light. Phenomenologically, both signs are observed, see Fig. 1. For example, the phononic is negative in SrTiO3 [5], Cu3TeO6 [6], and black phosphorus [24], as well as in all cuprates, whether undoped [3, 4], hole-doped [2], or electron-doped [21]. A positive phononic is observed in Fe2Mo3O8 [1], (Tb0.3Y0.7)2Ti2O7 [19], and RuCl[20, 39]. A remarkable feature of our findings is that changes sign to positive upon introducing 30% Y into TmVO4 (Fig. 5). This suggests that impurities, in much larger concentration in Tm0.7Y0.3VO4 relative to either TmVO4 or YVO4, is an important ingredient for the generation of a phonon thermal Hall effect.
It is instructive to look at the magnitude of . As argued by others [1, 6, 24], the relevant quantity for this is the ratio /, namely the Hall angle. In Fig. 6, we see that in our three samples / at K and T. This is quite typical. Indeed, the cuprate Mott insulators La2CuO4, Nd2CuO4, and Sr2CuO2Cl2 and the antiferromagnetic insulator Cu3TeO6 all have a Hall angle of about at the same temperature and field [6]. The Kitaev spin liquid candidate RuCl3 has a Hall angle of about . Note that the magnitude of varies by three orders of magnitude across those various materials (see Fig. 1).
It is worth noting that significantly larger thermal Hall angles have been observed in two cases so far: the Ce-doped cuprate Nd2CuO4 [21] and the Rh-doped iridate Sr2IrO4 [22]. In as-grown samples of Nd2CuO4 with 11% Ce and Sr2IrO4 with 5% Rh, / at K and T. Both materials are insulators with antiferromagnetic order at these dopings, and it was argued in the latter study that impurities embedded in an antiferromagnetic environment strongly promote the phonon thermal Hall effect [22], as suggested theoretically for a mechanism of resonant side-jump scattering of phonons [17].
V SUMMARY
We measured a non-zero thermal Hall conductivity in the vanadates TmVO4, YVO4 and Tm0.7Y0.3VO4. All evidence points to phonons as the heat carriers responsible for generating this Hall effect. The magnitude of the Hall response is similar in all three, with a Hall angle of / at K and T, comparable to the phonon Hall effect in several insulating materials. This shows that Tm3+ ions do not play an essential role in generating the Hall effect in these rare-earth vanadates. While the sign of in the two stoichiometric compounds TmVO4 and YVO4 is negative, we find a positive sign in the more disordered Tm0.7Y0.3VO4. This points to a special role played by impurities in this family of materials.
ACKNOWLEDGMENTS
We thank S. Fortier for his assistance with the experiments and A. Ataei, P. Fournier and I. Garate for stimulating discussions. L. T. acknowledges support from the Canadian Institute for Advanced Research (CIFAR) as a CIFAR Fellow and funding from the Institut Quantique, the Natural Sciences and Engineering Research Council of Canada (NSERC; PIN:123817), the Fonds de Recherche du Québec - Nature et Technologies (FRQNT), the Canada Foundation for Innovation (CFI), and a Canada Research Chair. This research was undertaken thanks in part to funding from the Canada First Research Excellence Fund.
Crystal growth and characterization performed at Stanford University was supported by the Air Force Office of Scientific Research under Award No. FA9550-20-1-0252. MPZ was also partially supported by a National Science Foundation Graduate Research Fellowship under Grant No. DGE-1656518.
APPENDIX
The thermal Hall effect can be measured in three ways and we have tested these methods on the hole-doped cuprate La2-xSrxCuO4 = 0.06 (LSCO).
(1) Steady-state method as function of temperature, also known as “-steps” (see Figs. 7 and 8). The magnetic field is kept fixed (at ) while the temperature of the sample is changed in discrete steps (2 to 3 K steps). At each temperature, the background voltage of the thermocouple that measures is recorded before the heat is applied to the sample. When the sample is in an equilibrium configuration, is measured.
The background voltage of the thermocouple is carefully subtracted from the measured heat-on signal from the thermocouple. Once the entire temperature range is covered, say from 10 K to 100 K, the field direction is reversed to -. The same procedure is followed for this field polarity. Then, the signal is antisymmetrized; , such that any symmetric contribution, coming from the longitudinal thermal gradient due to misalignment of the transverse contacts, is removed.
(2) Steady-state method as a function of field, also known as “-steps” (see Figs. 7 and 8). The temperature is kept fixed, and the magnetic field is changed in discrete steps (1 to 2 T steps). At a fixed temperature, a heat current is sent to the sample, and once it reaches equilibrium, the magnetic field is changed from to - in steps. For each field step , we define . Once the entire magnetic field range is covered, the temperature of the sample is changed and the same procedure is repeated.
(3) Field sweep method, also known as “-sweeps” (see Fig. 7). This method is similar to 2), but now the magnetic-field is continuously changed from to -.
We found satisfactory agreement between methods (1), (2), and (3), see Fig. 7.